# Reinforcement Learning in Vision: A Survey

### v2

@@ 3.1.4 Video Reasoning
Video reasoning extends the capabilities of MLLMs to process temporal dynamics, requiring not only spatial perception but also sequential understanding and causal inference. Recent works in this domain have proposed diverse approaches to tackle complex reasoning over video inputs. For instance, VQ-Insight (Zhang et al., 2025h) introduces a hierarchical reward design and self-consistency voting mechanism tailored to the question–answering process over long videos. TW-GRPO (Dang et al., 2025) combines token wise credit assignment with GRPO-style optimization to improve fine-grained temporal alignment between textual reasoning and video evidence. Meanwhile, several R1-style frameworks have been developed to unlock video understanding in complex real-world or egocentric settings. EgoVLM (Vinod et al., 2025) and VAU-R1 (Zhu et al., 2025b) focus on egocentric video reasoning with visual memory and utility-based rewards. DeepVideo-R (Park et al., 2025a) integrates dense video encodings and external reward functions to supervise long-horizon reasoning. TimeMaster (Zhang et al., 2025e) explicitly structures temporal abstraction and reasoning via curriculum learning, while VideoR1 (Feng et al., 2025) proposes a scalable RL framework for video-based QA tasks across multiple domains. Collectively, these works highlight the importance of aligning temporal representations with language trajectories through reinforcement learning, paving the way for robust and generalizable video reasoning agents.
@@ 3.2 Visual Generation
Figure 5: Three reward paradigms for RL-based image generation. (a) Human-Centric Preference Optimization: aligns outputs with human aesthetic scores (HPS (Wu et al., 2023b), ImageReward (Xu et al., 2023)); (b) Multimodal Reasoning-Based Evaluation: scores images via multimodal reasoning consistency (UnifiedReward (Wang et al., 2025j), PARM (Guo et al., 2025d)); (c) Metric-Driven Objective Optimization: minimizes task-specific quantitative metrics such as FID and IoU.
@@ 3.2.1 Image Generation
RL for image generation models departs from the language counterpart in both action space continuous diffusion steps or prompt refinements and reward design, which must capture perceptual quality, text–image alignment, and subject fidelity. A first family learns an explicit visual reward model: ImageReward (Xu et al., 2023) supplies human-preference scores that drive policy-gradient fine-tuning of diffusion backbones in DiffPPO (Xiao et al., 2024), Dpok (Fan et al., 2023a), and FocusDiff (Pan et al., 2025). A second line bypasses reward modelling by optimising pairwise or unary preferences: DDPO (Black et al., 2023), DiffusionDPO (Wallace et al., 2024), Diffusion-KTO (Li et al., 2024b), and DiffusionRPO (Gu et al., 2024) treat denoising trajectories as MDPs and apply R1/GRPO updates to maximise comparative human feedback. Beyond alignment, works such as PrefPaint (Liu et al., 2024a), Parrot (Lee et al., 2024), and RLRF (Rodriguez et al., 2025b) craft multi-objective or render-and-compare rewards to refine aesthetics, diversity, or vector graphics. RL has also been used to inject reasoning and prompt adaptation: ReasonGen-R1 (Zhang et al., 2025i), GoT-R1 (Duan et al., 2025) and RePrompt (Wu et al., 2025b) first generate textual plans or improved prompts, then reinforce the generator for coherent scene synthesis. Finally, personalisation methods—DPG-T2I (Wei et al., 2024), RPO (Miao et al., 2024a), and B2-DiffuRL (Hu et al., 2025a) optimize rewards that measure subject fidelity under scarce reference images. Collectively, these studies show that RL, armed with perceptual or preference-based rewards, can steer diffusion models toward higher realism, stronger prompt faithfulness, controllable layout, and user-specific appearance capabilities difficult to achieve with likelihood training alone.
@@ 3.2.2 Video Generation
Applying RL to video generation introduces challenges absent in images: rewards must capture temporal coherence, motion naturalness, and text–video alignment across hundreds of frames. Early work such as InstructVideo (Yuan et al., 2024) repurposed image scorers and applied PPO to refine short clips, whereas VideoRM (Wu et al., 2024) and VideoReward (Liu et al., 2025b) learn dedicated preference models that grade entire sequences for smoothness, consistency and prompt faithfulness. Building on GRPO/R1, DanceGRPO (Xue et al., 2025) shows that group-normalized returns stabilize long-horizon optimization and boost aesthetic and alignment scores across diffusion and flow backbones.
Beyond preference alignment, specialized rewards tackle domain-specific goals. GAPO (Zhu et al., 2025a) exploits gap-aware ranking to fine-tune anime videos; Phys-AR (Lin et al., 2025) penalizes violations of physics to yield plausible trajectories; and InfLVG (Fang et al., 2025) trains an inference-time policy that retains only context tokens beneficial for multi-scene coherence. Auxiliary critics further facilitate training: VideoScore (He et al., 2024b) and Gradeo (Mou et al., 2025) offer explainable, multi-factor scores, while TeViR (Chen et al., 2025d) feeds imagined video roll-outs as dense rewards to downstream control agents. Collectively, these studies demonstrate that carefully crafted sequence level rewards and group-relative policy updates are key to producing temporally consistent, semantically aligned, and physically plausible videos capabilities difficult to obtain with likelihood training alone.
@@
RL for 3D generation differs from 2D and video tasks as rewards operate on volumetric structures or rendered views, often requiring expensive render-and-compare loops. DreamCS (Zou et al., 2025) pioneers this paradigm by framing text-to-mesh synthesis as a Markov decision process: a diffusion backbone proposes coarse shapes, then a policy refines vertex positions under a reward that jointly measures silhouette IoU, CLIP text-mesh alignment, and mesh smoothness. On the reward side, DreamReward (Ye et al., 2024) introduces a large-scale human preference dataset of rendered 3-D assets and trains a geometry-aware critic that scores prompts, renders, and latent SDFs; the learned signal enables stable PPO fine-tuning of existing text-to-shape models.
A complementary line adopts direct preference optimization. DreamDPO (Zhou et al., 2025c) extends DPO to NeRF and mesh diffusion models by sampling paired 3D outputs and maximizing the margin dictated by human comparisons, achieving superior prompt fidelity without an explicit value network. Finally, Nabla-R2D3 (Liu et al., 2025d) tackles interactive scene editing: the agent sequentially adds, deletes, or transforms objects in a 3D scene; reward is computed via real-time rasterized views and task-specific validators (e.g., occupancy, spatial relations). Group-relative policy optimization (R2D3) stabilizes training despite sparse feedback. Together, these studies demonstrate that RL, equipped with geometry-aware or render-based rewards, provides an effective mechanism for controlling structural integrity, text alignment, and interactive editability capabilities that standard likelihood or score-distillation objectives struggle to capture in the 3D domain.
@@ 3.3.1 Unified RL
UniRL (Mao et al., 2025) exemplifies this paradigm: a visual autoregressive backbone is first instruction-tuned, then jointly fine-tuned on VQA, captioning and image generation with a blended reward measuring textual correctness, CLIP-based alignment, and aesthetic quality. CoRL (Jiang et al., 2025b) pushes the idea further by alternating “co-understanding” and “co-generation” batches within the same GRPO step. To address inefficiency in dense token spaces, SelfTok (Wang et al., 2025a) discretises multi-modal actions into a self-evolving token set and demonstrates that a single RL head can govern retrieval, grounding, and synthesis with minimal extra parameters. Finally, HermesFlow (Yang et al., 2025c) couples an autoregressive text module with a rectified flow image decoder under one cross-task reward, illustrating that diffusion-style and language-style policies can be harmonized through unified reinforcement updates. Together, these works suggest that sharing a common RL objective across tasks not only reduces training cost but also encourages emergent cross modal generalization unavailable to isolated, task specific fine-tuning.
@@ 3.3.2 Task specific RL
In contrast to the unified approaches of §3.3.1, task-specific RL confines the reward signal to a single downstream objective, optimizing one functional head while leaving other capabilities untouched. VARGPT-v1.1 (Zhuang et al., 2025) exemplifies this strategy: although the underlying visual autoregressive model can handle both understanding and generation, its RL stage targets only visual generation with DPO. Similarly, Emu3 (Wang et al., 2024b) introduces RL exclusively to polish its image generation branch, which leveraging pair wise human preferences. For the multimodal understanding abilities of model (e.g.,  captioning, VQA), the work just train this part by task specific fine-tuning alone.
@@ 3.4.1 GUI Automation
Modern GUI RL research frames screen understanding and action prediction as a vision–language decision process, then employs rule-based or preference rewards to close the perception–action loop. On desktop and web interfaces, GUI-R1 (Luo et al., 2025) introduces an R1-style rule set that maps click success, text entry, and scroll validity to dense rewards. UI-R1 (Lu et al., 2025) adds GRPO with a novel action-specific KL term to stabilize long-horizon plans, while SE-GUI (Yuan et al., 2025b) applies self-evolutionary filtering to distil high-fidelity trajectories. Focusing on trajectory reuse, UIShift (Gao et al., 2025b) formulates an inverse dynamics objective that lets MLLM learn actions from unlabeled GUI pairs and then refines them via RL. Complementary preference-based frameworks include LPO (Tang et al., 2025) that rewards spatial proximity for precise clicks. ProgRM (Zhang et al., 2025b) injects program-level logical checks, and RUIG (Zhang et al., 2023) leverages instruction grounding with reinforcement signals. Tool-specific baselines such as Ui-tars (Qin et al., 2025) offer larger action vocabularies yet still rely on rule-driven RL for robust execution.
Mobile scenarios introduce latency and on-device constraints. AgentCPM-GUI (Zhang et al., 2025j) compresses the action space and conducts GRPO fine-tuning. MobileGUI-RL (Shi et al., 2025) advances this line via online RL with task-level rewards to improve exploration under limited memory, and Mobile-R1 (Gu et al., 2025) extends interactive multi-turn RL to correct error cascades during long tasks. At inference, GTA1 (Yang et al., 2025e) samples multiple action candidates and employs a judge model to pick the best, effectively trading compute for higher success rates. Additional, light-weight models such as Appvlm (Papoudakis et al., 2025) demonstrate that modest-sized MLLM, after GRPO fine-tuning, can control smartphone apps with competitive precision. To adaptively reduce the thinking length, the TON (Wang et al., 2025b) proposes a thought-dropout solution during supervised fine-tuning stage, then GRPO skill adaptivley skip unnecessary reasoning process for efficiently thinking.
@@ 3.4.2 Visual Navigation
RL-driven visual navigation research now couples large vision–language models with embodied control, employing group-normalized or time-decayed returns to maintain long-horizon stability. OctoNav-R1 (Gao et al., 2025a) exploits a hybrid RL pipeline with a “think-before-action” ability for VLA model, then translate egocentric frames into low-level actions. Focusing on dataset efficiency, VLN-R1 (Qi et al., 2025) builds an end-to-end navigator and introduces a time-decayed reward to handle continuous trajectories. At the system level, Flare (Hu et al., 2024a) demonstrates that fine-tuning a multi-task robotics policy with large-scale RL in simulation can generalize to real-world household tasks. Complementary advances include More (Zhao et al., 2025), which augments omni-directional inputs with memory-guided policy distillation, and RAPID (Kim et al., 2025), which integrates pose priors for faster convergence in unseen layouts. These works show that using temporal rewards, memory sampling, and environment priors with GRPO/PPO helps VLA agents navigate more reliably and efficiently.
@@ 3.4.3 Visual Manipulation
Visual manipulation tasks (e.g.,  object relocation, tool use, and multi-step rearrangement) require fine-grained perception and long-horizon planning. Recent works (Chen et al., 2025g; Shu et al., 2025) integrate reinforcement learning with vision–language models to enhance generalization, interactivity, and policy consistency. TGRPO (Chen et al., 2025g) introduces a task-grounded reward formulation and group normalized updates to stabilize training for open-ended object manipulation. RFTF (Shu et al., 2025) applies rule-based rewards to support interactive table top tasks and emphasizes training with minimal human supervision. Meanwhile, RLVLA (Liu et al., 2025c) and VLA-RL (Lu et al., 2025b) explore curriculum-based or progressive reinforcement learning for VLM-based robot agents, achieving high success rates across diverse rearrangement environments.
Building on this, ConRFT (Chen et al., 2025e) and iRe-VLA (Guo et al., 2025b) introduce consistency-aware and instruction-refinement strategies respectively, using RL to align visual predictions with physical interaction outcomes. RIPT-VLA focuses on interactive prompting during manipulation, bridging LLM planning and low-level control through reinforced feedback (Tan et al., 2025b). Finally, ReinBot (Zhang et al., 2025c) leverages multimodal rollouts and preference-based updates to improve real-world manipulation robustness (Zhang et al., 2025c). Collectively, these studies highlight the role of vision-language reasoning, structured reward design, and RL-based refinement in advancing embodied manipulation under complex, language-conditioned settings.
@@
Table 2: Overview of evaluation metrics in large-model visual RL. Each task family is broken down into Task Metrics (RL-free external benchmarks), Reward Metrics (how the learning signal is computed), and Model-state Metrics (diagnostics tracked during optimization).
@@
Multimodal LLMs/ VLMs Comprehensive Evaluation: e.g.,  MME (Fu et al., 2023), SEED-Bench (Li et al., 2024a), VQA v2 (Goyal et al., 2017), MMBench (Liu et al., 2024b).
OCR: e.g.,  TextVQA (Singh et al., 2019), OCR-VQA (Mishra et al., 2019), OCRBench (Liu et al., 2023b).
Mathematical: e.g.,  MathVista (Lu et al., 2023).
Documentation: e.g.,  ChartQA (Lu et al., 2023), DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022).
Multilingual: e.g.,  CMMMU (Zhang et al., 2024a), CMMU (He et al., 2024c).
Reward from Human Preference: e.g.,  InstructGPT (Ouyang et al., 2022), LLaVA-RLHF (Sun et al., 2023).
Verifiable Rewards: e.g.,  Deepseekmath (Shao et al., 2024b), IoU (Rezatofighi et al., 2019), CodeRL (Le et al., 2022).
Reward from Model Preferencee.g., CriticGPT (McAleese et al., 2024).
Output length monitoring: e.g.,  DPO (Rafailov et al., 2023).
KL-divergence: e.g.,  InstructGPT (Ouyang et al., 2022).
@@
Image Fidelity & Diversity: e.g.,  FID (Heusel et al., 2017), Inception Score (IS) (Salimans et al., 2016).
Pixel-level Reconstruction: e.g.,  PSNR (Wang et al., 2004), SSIM (Wang et al., 2004).
Semantic Alignment: e.g.,  Geneval (Ghosh et al., 2023), T2I-CompBench (Huang et al., 2023), CLIP Score (Radford et al., 2021), Frechet CLIP Distance (Betzalel et al., 2022).
Video Fidelity & Diversity: e.g.,  FVD (Unterthiner et al., 2019), Video IS (Saito et al., 2020).
Reward from Human Preferencee.g., ImageReward (Xu et al., 2023), HPS (Wu et al., 2023b), HPS V2 (Wu et al., 2023b), Pick-a-pic (Kirstain et al., 2023), VideoReward (Liu et al., 2025b), RichHF-18K (Liang et al., 2024).
Verifiable Rewardse.g., MotionPrompt (Nam et al., 2025), DSPO (Cai et al., 2025b), Instructrl4pix (Li et al., 2024c).
Reward from Model Preferencee.g., VideoPrefer (Wu et al., 2024), PARM (Zhang et al., 2025f).
Denoising Trajectory Diagnostics: e.g.,  VARD (Dai et al., 2025), Inversion-DPO (Li et al., 2025f).
KL-divergence: e.g.,  DDPO (Black et al., 2023), VARD (Dai et al., 2025).
Unified Models Generation–oriented Task Metricse.g., GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024b), ParaPrompts (Wu et al., 2025d).
Understanding–oriented Task Metricse.g., MME (Fu et al., 2023), POPE (Li et al., 2023).
Unified Rewardse.g., UniRL (Mao et al., 2025), CoRL (Jiang et al., 2025b).
Task-specific Rewardse.g., Vargpt-v1. 1 (Zhuang et al., 2025), Emu3 (Wang et al., 2024b).
Generation–Understanding Imbalancee.g., UniRL (Mao et al., 2025).
KL-divergencee.g., HermesFlow (Yang et al., 2025c).
@@
GUI Action Accuracy (offline)e.g., ScreenSpot (Cheng et al., 2024), ScreenSpot-Pro (Li et al., 2025c), Ui-vision (Nayak et al., 2025).
Task Success rate (online): e.g., Webarena (Zhou et al., 2023), Osworld (Xie et al., 2024), Windows agent arena (Bonatti et al., 2024).
Model-based Evaluation: e.g., Agentrewardbench (Lù et al., 2025), Webworld (Chae et al., 2024), Digirl (Bai et al., 2024).
Reward from Human Preferencee.g., RFTF (Shu et al., 2025).
Rule-based Success (binary) (IoU, Action Accuracy): e.g., UI-R1 (Lu et al., 2025), ARPO (Lu et al., 2025a), VLA-RL (Lu et al., 2025b), iRe-VLA (Guo et al., 2025b).
Model Preference Critic: e.g., ProgRM (Zhang et al., 2025b).
Dense Shaping (distance or coverage): e.g., LPO (Tang et al., 2025), Gui-r1 (Luo et al., 2025).
Action trajectory length: e.g., Osworld (Xie et al., 2024).
KL penalty for policy stabilitye.g., GUI-R1 (Luo et al., 2025), UI-R1 (Lu et al., 2025).
Output length monitoring: e.g., UI-R1 (Lu et al., 2025).

### v3

@@ 3.1.4 Video Reasoning
Video reasoning extends the capabilities of MLLMs to process temporal dynamics, requiring not only spatial perception but also sequential understanding and causal inference. Recent works in this domain have proposed diverse approaches to tackle complex reasoning over video inputs. For instance, VQ-Insight [zhang2025vq] introduces a hierarchical reward design and self-consistency voting mechanism tailored to the question–answering process over long videos. TW-GRPO [dang2025reinforcing] combines token wise credit assignment with GRPO-style optimization to improve fine-grained temporal alignment between textual reasoning and video evidence. Meanwhile, several R1-style frameworks have been developed to unlock video understanding in complex real-world or egocentric settings. EgoVLM [vinod2025egovlm] and VAU-R1 [zhu2025vau] focus on egocentric video reasoning with visual memory and utility-based rewards. DeepVideo-R [park2025deepvideo] integrates dense video encodings and external reward functions to supervise long-horizon reasoning. TimeMaster [zhang2025timemaster] explicitly structures temporal abstraction and reasoning via curriculum learning, while VideoR1 [videor1] proposes a scalable RL framework for video-based QA tasks across multiple domains. Collectively, these works highlight the importance of aligning temporal representations with language trajectories through reinforcement learning, paving the way for robust and generalizable video reasoning agents.
@@ 3.2.1 Image Generation
RL for image generation models departs from the language counterpart in both action space continuous diffusion steps or prompt refinements and reward design, which must capture perceptual quality, text–image alignment, and subject fidelity. A first family learns an explicit visual reward model: ImageReward [xu2023imagereward] supplies human-preference scores that drive policy-gradient fine-tuning of diffusion backbones in DiffPPO [xiao2024diffppo], Dpok [fan2023dpok], and FocusDiff [pan2025focusdiff]. A second line bypasses reward modelling by optimising pairwise or unary preferences: DDPO [black2023ddpo], DiffusionDPO [wallace2024diffusion], Diffusion-KTO [li2024aligning], and DiffusionRPO [gu2024diffusion] treat denoising trajectories as MDPs and apply R1/GRPO updates to maximise comparative human feedback. Beyond alignment, works such as PrefPaint [liu2024prefpaint], Parrot [lee2024parrot], and RLRF [rodriguez2025rendering] craft multi-objective or render-and-compare rewards to refine aesthetics, diversity, or vector graphics. RL has also been used to inject reasoning and prompt adaptation: ReasonGen-R1 [zhang2025reasongen], GoT-R1 [duan2025got] and RePrompt [wu2025reprompt] first generate textual plans or improved prompts, then reinforce the generator for coherent scene synthesis. Finally, personalisation methods—DPG-T2I [wei2024powerful], RPO [miao2024subject], and B2-DiffuRL [hu2025towards] optimize rewards that measure subject fidelity under scarce reference images. Collectively, these studies show that RL, armed with perceptual or preference-based rewards, can steer diffusion models toward higher realism, stronger prompt faithfulness, controllable layout, and user-specific appearance capabilities difficult to achieve with likelihood training alone.
@@ 3.2.2 Video Generation
Applying RL to video generation introduces challenges absent in images: rewards must capture temporal coherence, motion naturalness, and text–video alignment across hundreds of frames. Early work such as InstructVideo [yuan2024instructvideo] repurposed image scorers and applied PPO to refine short clips, whereas VideoRM [wu2024boosting] and VideoReward [liu2025improving] learn dedicated preference models that grade entire sequences for smoothness, consistency and prompt faithfulness. Building on GRPO/R1, DanceGRPO [xue2025dancegrpo] shows that group-normalized returns stabilize long-horizon optimization and boost aesthetic and alignment scores across diffusion and flow backbones.
Beyond preference alignment, specialized rewards tackle domain-specific goals. GAPO [zhu2025aligning] exploits gap-aware ranking to fine-tune anime videos; Phys-AR [lin2025reasoning] penalizes violations of physics to yield plausible trajectories; and InfLVG [fang2025inflvg] trains an inference-time policy that retains only context tokens beneficial for multi-scene coherence. Auxiliary critics further facilitate training: VideoScore [he2024videoscore] and Gradeo [mou2025gradeo] offer explainable, multi-factor scores, while TeViR [chen2025tevir] feeds imagined video roll-outs as dense rewards to downstream control agents. Collectively, these studies demonstrate that carefully crafted sequence level rewards and group-relative policy updates are key to producing temporally consistent, semantically aligned, and physically plausible videos capabilities difficult to obtain with likelihood training alone.
@@
RL for 3D generation differs from 2D and video tasks as rewards operate on volumetric structures or rendered views, often requiring expensive render-and-compare loops. DreamCS [zou2025dreamcs] pioneers this paradigm by framing text-to-mesh synthesis as a Markov decision process: a diffusion backbone proposes coarse shapes, then a policy refines vertex positions under a reward that jointly measures silhouette IoU, CLIP text‐mesh alignment, and mesh smoothness. On the reward side, DreamReward [ye2024dreamreward] introduces a large‐scale human preference dataset of rendered 3-D assets and trains a geometry-aware critic that scores prompts, renders, and latent SDFs; the learned signal enables stable PPO fine-tuning of existing text-to-shape models.
A complementary line adopts direct preference optimization. DreamDPO [zhou2025dreamdpo] extends DPO to NeRF and mesh diffusion models by sampling paired 3D outputs and maximizing the margin dictated by human comparisons, achieving superior prompt fidelity without an explicit value network. Finally, Nabla-R2D3 [liu2025nabla] tackles interactive scene editing: the agent sequentially adds, deletes, or transforms objects in a 3D scene; reward is computed via real-time rasterized views and task-specific validators (e.g., occupancy, spatial relations). Group-relative policy optimization (R2D3) stabilizes training despite sparse feedback. Together, these studies demonstrate that RL, equipped with geometry-aware or render-based rewards, provides an effective mechanism for controlling structural integrity, text alignment, and interactive editability capabilities that standard likelihood or score-distillation objectives struggle to capture in the 3D domain.
Figure 5: Three reward paradigms for RL-based image generation. (a) Human-Centric Preference Optimization: aligns outputs with human aesthetic scores (HPS [wu2023human], ImageReward [xu2023imagereward]); (b) Multimodal Reasoning-Based Evaluation: scores images via multimodal reasoning consistency (UnifiedReward [wang2025unified], PARM [guo2025can]); (c) Metric-Driven Objective Optimization: minimizes task-specific quantitative metrics such as FID and IoU.
3.2.4 Reward Design and Preference Modeling for Visual Generation
As shown in Figure 5, the reward design in RL-based visual generation can broadly be divided into three paradigms: (a) Human-Centric Preference Optimization (using human feedback or learned reward models for video alignment and fidelity). Representative works include ImageReward [xu2023imagereward], which pioneered large-scale text-to-image reward modeling from human comparisons and provided a critic for diffusion fine-tuning. Dpok [fan2023dpok] and DiffPPO [xiao2024diffppo] adapted PPO with KL regularization and human aesthetic scores to enhance alignment and fidelity. Extending these ideas to video, VideoRM [wu2024boosting] and VideoReward [liu2025improving] trained reward models on large-scale human (and GPT-assisted) video comparisons, enabling RL fine-tuning of text-to-video generation. (b) Multimodal Reasoning-Based Evaluation (using a question-answering or reasoning module to score consistency between videos and textual descriptions). For example, LLaVA-Reward [zhou2025multimodalllmscustomizedreward] leverages a large VLM to evaluate images across multiple dimensions including alignment, fidelity, and safety, providing a multi-faceted reward signal that better approximates human judgment. and (c) Metric-Driven Objective Optimization (directly optimizing quantitative metrics like FID or IoU for quality).
@@ 3.3.1 Unified RL
UniRL [mao2025unirl] exemplifies this paradigm: a visual autoregressive backbone is first instruction-tuned, then jointly fine-tuned on VQA, captioning and image generation with a blended reward measuring textual correctness, CLIP-based alignment, and aesthetic quality. CoRL [jiang2025co] pushes the idea further by alternating “co-understanding” and “co-generation” batches within the same GRPO step. To address inefficiency in dense token spaces, SelfTok [wang2025discrete] discretises multi-modal actions into a self-evolving token set and demonstrates that a single RL head can govern retrieval, grounding, and synthesis with minimal extra parameters. Finally, HermesFlow [yang2025hermesflow] couples an autoregressive text module with a rectified flow image decoder under one cross-task reward, illustrating that diffusion-style and language-style policies can be harmonized through unified reinforcement updates. Together, these works suggest that sharing a common RL objective across tasks not only reduces training cost but also encourages emergent cross modal generalization unavailable to isolated, task specific fine-tuning.
@@ 3.3.2 Task specific RL
In contrast to the unified approaches of §3.3.1, task-specific RL confines the reward signal to a single downstream objective, optimizing one functional head while leaving other capabilities untouched. VARGPT-v1.1 [zhuang2025vargpt] exemplifies this strategy: although the underlying visual autoregressive model can handle both understanding and generation, its RL stage targets only visual generation with DPO. Similarly, Emu3 [wang2024emu3] introduces RL exclusively to polish its image generation branch, which leveraging pair wise human preferences. For the multimodal understanding abilities of model (e.g.,  captioning, VQA), the work just train this part by task specific fine-tuning alone.
@@ 3.4.1 GUI Automation
Modern GUI RL research frames screen understanding and action prediction as a vision–language decision process, then employs rule-based or preference rewards to close the perception–action loop. On desktop and web interfaces, GUI-R1 [luo2025gui] introduces an R1-style rule set that maps click success, text entry, and scroll validity to dense rewards. UI-R1 [lu2025ui] adds GRPO with a novel action-specific KL term to stabilize long-horizon plans, while SE-GUI [yuan2025enhancing] applies self-evolutionary filtering to distil high-fidelity trajectories. Focusing on trajectory reuse, UIShift [gao2025uishift] formulates an inverse dynamics objective that lets MLLM learn actions from unlabeled GUI pairs and then refines them via RL. Complementary preference-based frameworks include LPO [tang2025lpo] that rewards spatial proximity for precise clicks. ProgRM [zhang2025progrm] injects program-level logical checks, and RUIG [zhang2023reinforced] leverages instruction grounding with reinforcement signals. Tool-specific baselines such as Ui-tars [qin2025ui] offer larger action vocabularies yet still rely on rule-driven RL for robust execution.
Mobile scenarios introduce latency and on-device constraints. AgentCPM-GUI [zhang2025agentcpm] compresses the action space and conducts GRPO fine-tuning. MobileGUI-RL [shi2025mobilegui] advances this line via online RL with task-level rewards to improve exploration under limited memory, and Mobile-R1 [gu2025mobile] extends interactive multi-turn RL to correct error cascades during long tasks. At inference, GTA1 [yang2025gta1] samples multiple action candidates and employs a judge model to pick the best, effectively trading compute for higher success rates. Additional, light-weight models such as Appvlm [papoudakis2025appvlm] demonstrate that modest-sized MLLM, after GRPO fine-tuning, can control smartphone apps with competitive precision. To adaptively reduce the thinking length, the TON [ton] proposes a thought-dropout solution during supervised fine-tuning stage, then GRPO skill adaptivley skip unnecessary reasoning process for efficiently thinking.
@@ 3.4.2 Visual Navigation
RL‐driven visual navigation research now couples large vision–language models with embodied control, employing group‐normalized or time‐decayed returns to maintain long‐horizon stability. OctoNav-R1 [gao2025octonav] exploits a hybrid RL pipeline with a “think-before-action” ability for VLA model, then translate egocentric frames into low-level actions. Focusing on dataset efficiency, VLN-R1 [qi2025vln] builds an end-to-end navigator and introduces a time-decayed reward to handle continuous trajectories. At the system level, Flare [hu2024flare] demonstrates that fine-tuning a multi-task robotics policy with large-scale RL in simulation can generalize to real-world household tasks. Complementary advances include More [zhao2025more], which augments omni-directional inputs with memory-guided policy distillation, and RAPID [kim2025rapid], which integrates pose priors for faster convergence in unseen layouts. These works show that using temporal rewards, memory sampling, and environment priors with GRPO/PPO helps VLA agents navigate more reliably and efficiently.
@@ 3.4.3 Visual Manipulation
Visual manipulation tasks (e.g.,  object relocation, tool use, and multi-step rearrangement) require fine-grained perception and long-horizon planning. Recent works [chen2025tgrpo, shu2025rftf] integrate reinforcement learning with vision–language models to enhance generalization, interactivity, and policy consistency. TGRPO [chen2025tgrpo] introduces a task-grounded reward formulation and group normalized updates to stabilize training for open-ended object manipulation. RFTF [shu2025rftf] applies rule-based rewards to support interactive table top tasks and emphasizes training with minimal human supervision. Meanwhile, RLVLA [liu2025can] and VLA-RL [lu2025vla] explore curriculum-based or progressive reinforcement learning for VLM-based robot agents, achieving high success rates across diverse rearrangement environments.
Building on this, ConRFT [chen2025conrft] and iRe-VLA [guo2025improving] introduce consistency-aware and instruction-refinement strategies respectively, using RL to align visual predictions with physical interaction outcomes. RIPT-VLA focuses on interactive prompting during manipulation, bridging LLM planning and low-level control through reinforced feedback [tan2025interactive]. Finally, ReinBot [zhang2025reinbot] leverages multimodal rollouts and preference-based updates to improve real-world manipulation robustness [zhang2025reinbot]. Collectively, these studies highlight the role of vision-language reasoning, structured reward design, and RL-based refinement in advancing embodied manipulation under complex, language-conditioned settings.
@@
TABLE II: Overview of evaluation metrics in large-model visual RL. Each task family is broken down into Task Metrics (RL-free external benchmarks), Reward Metrics (how the learning signal is computed), and Model-state Metrics (diagnostics tracked during optimization).
@@
Multimodal LLMs/ VLMs Comprehensive Evaluation: e.g.,  MME [fu2023mme], SEED-Bench [li2024seed], VQA v2 [goyal2017making], MMBench [liu2024mmbench].
OCR: e.g.,  TextVQA [singh2019towards], OCR-VQA [mishra2019ocr], OCRBench [liu2023hidden].
Mathematical: e.g.,  MathVista [lu2023mathvista].
Documentation: e.g.,  ChartQA [lu2023mathvista], DocVQA [mathew2021docvqa], InfoVQA [mathew2022infographicvqa].
Multilingual: e.g.,  CMMMU [zhang2024cmmmu], CMMU [he2024cmmu].
[…]
Reward from Human Preference: e.g.,  InstructGPT [ouyang2022training], LLaVA-RLHF [sun2023aligning].
Verifiable Rewards: Deepseekmath [shao2024deepseekmath], IoU [rezatofighi2019generalized], CodeRL [le2022coderl].
Reward from Model Preferencee.g., CriticGPT [mcaleese2024llm].
[…]
Output length monitoring: e.g.,  DPO [rafailov2023direct].
KL-divergence: e.g.,  InstructGPT [ouyang2022training].
[…]
@@
Image Fidelity & Diversity: e.g.,  FID [heusel2017gans], Inception Score (IS) [salimans2016improved].
Pixel-level Reconstruction: e.g.,  PSNR [wang2004image], SSIM [wang2004image].
Semantic Alignment: e.g.,  Geneval [ghosh2023geneval], T2I-CompBench [huang2023t2i], CLIP Score [radford2021learning], Frechet CLIP Distance [betzalel2022study].
Video Fidelity & Diversity: e.g.,  FVD [unterthiner2019fvd], Video IS [saito2020train].
[…]
Reward from Human Preferencee.g., ImageReward [xu2023imagereward], HPS [wu2023human], HPS V2 [wu2023human], Pick-a-pic [kirstain2023pick], VideoReward [liu2025improving], RichHF-18K [liang2024rich].
Verifiable Rewardse.g., MotionPrompt [nam2025optical], DSPO [cai2025dspo], Instructrl4pix [li2024instructrl4pix].
Reward from Model Preferencee.g., VideoPrefer [wu2024boosting], PARM [zhang2025let].
[…]
Denoising Trajectory Diagnostics: e.g.,  VARD [dai2025vard], Inversion-DPO [li2025inversion].
KL-divergence: e.g.,  DDPO [black2023ddpo], VARD [dai2025vard].
[…]
Unified Models Generation–oriented Task Metricse.g., GenEval [ghosh2023geneval], DPG-Bench [hu2024ella], ParaPrompts [wu2025paragraph].
Understanding–oriented Task Metricse.g., MME [fu2023mme], POPE [li2023evaluating].
[…]
Unified Rewardse.g., UniRL [mao2025unirl], CoRL [jiang2025co].
Task-specific Rewardse.g., Vargpt-v1. 1 [zhuang2025vargpt], Emu3 [wang2024emu3].
[…]
Generation–Understanding Imbalancee.g., UniRL [mao2025unirl].
KL-divergencee.g., HermesFlow [yang2025hermesflow]
[…]
@@
GUI Action Accuracy (offline)e.g., ScreenSpot [seeclick], ScreenSpot-Pro [screenspotpro], Ui-vision [uivision].
Task Success rate (online): e.g., Webarena [webarena], Osworld [osworld], Windows agent arena [windowarena].
Model-based Evaluation: e.g., Agentrewardbench [agentrewardbench], Webworld [webworld], Digirl [digirl].
Reward from Human Preferencee.g., RFTF [shu2025rftf].
Navigation Success & SPL:
[…]
Rule-based Success (binary) (IoU, Action Accuracy): e.g., UI-R1 [lu2025ui], ARPO [arpo], VLA-RL [lu2025vla].
Model Preference Critic: e.g., ProgRM [zhang2025progrm].
Dense Shaping (distance or coverage): e.g., LPO [tang2025lpo], Gui-r1 [luo2025gui].
[…]
Action trajectory length: e.g., Osworld [osworld].
KL penalty for policy stabilitye.g., GUI-R1 [luo2025gui], UI-R1 [lu2025ui].
[…]